Solve the systems using substitution.
1) 3x - 11y = 16
x + y = 3
2) 4x - y = 4
x + 2y = 10
HELP I DONT KNOW WHAT IM DOING

Answer by grishma.kshatriya(27)

Solution :

The substitution method is used to eliminate one of the variables by replacement when solving a system of equations.
(1)
3x - 11y = 16..........(1)
x + y = 3..............(2)
We can write the (2) equation as x = 3 - y
Substituting the value x in equation (1) we get ,
3x - 11y = 16
3(3 - y ) - 11y = 16
9 - 3y - 11y = 16
9 - 14y = 16
-14y = 16 - 9
-14y = 7
So y = -1/2
Replug the y=-1/2 in (2) we get ,
x - 1/2 = 3
x = 3+1/2
x = 7/2
So we have the value as x = 7/2 and y = -1/2

(2)
4x - y = 4.......(1)
x + 2y = 10......(2)
from (2) x = 10 - 2y
Plug it in (1)
4(10 - 2y) - y = 4
40 - 8y - y = 4
40 - 9y = 4
-9y = -40 + 4
-9y = -36
y = 4
Replug in equation (2)
x+ 2y = 10
x + 2(4) = 10
x = 10 - 8
x = 2

So we have the value as x = 2 and y = 4
Thank you.